a cross section through a very large no conducting slab of thikness d=9.40 mm
 and uniformvolume charge density p = 5.80fC/m3 .the origin of an x axis is at centre.what is the magnitude of the slab electric field atan x cordinate of a)0,b)2mm,c)4.70 and d)26 mm?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Volume charge density = 5.80 fC/m3
Thickness d = 9.40 mm
At 0, Electric field will be zero.
At 2 mm
The Electric field will be  
E.F. = volume density * volume
=  5.80 * 18.8
= 109.04
At 4.70 mm
Electric Field will be
E.F. = volume density* volume
= 5.80* 9.40
= 256.244
At 26mm
E.F. = volume density * volume
= 5.80 * 244.4
= 1417.52

Regards
Arun (askIITians forum expert)